### v1

# DuSEGO: Dual Second-order Equivariant Graph Ordinary Differential Equation

Abstract.
Graph Neural Networks (GNNs) with equivariant properties have achieved significant success in modeling complex dynamic systems and molecular properties. However, their expressiveness ability is limited by: (1) Existing methods often overlook the over-smoothing issue caused by traditional GNN models, as well as the gradient explosion or vanishing problems in deep GNNs. (2) Most models operate on first-order information, neglecting that the real world often consists of second-order systems, which further limits the model’s representation capabilities. To address these issues, we propose the Dual Second-order Equivariant Graph Ordinary Differential Equation (DuSEGO) for equivariant representation. Specifically, DuSEGO apply the dual second-order equivariant graph ordinary differential equations (Graph ODEs) on graph embeddings and node coordinates, simultaneously. Theoretically, we first prove that DuSEGO maintains the equivariant property. Furthermore, we provide theoretical insights showing that DuSEGO effectively alleviates the over-smoothing problem in both feature representation and coordinate update. Additionally, we demonstrate that the proposed DuSEGO mitigates the exploding and vanishing gradients problem, facilitating the training of deep multi-layer GNNs. Extensive experiments on benchmark datasets validate the superiority of the proposed DuSEGO compared to baselines.
Graph Ordinary Differential Equation, Graph Neural Network, Equivariant Graph
Co-corresponding authors.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: TKDDconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Neural networks
1. Introduction
GNNs (kipf2017semisupervised, ; Yao2023ImprovingTE, ; yin2024dynamic, ; yin2023coco, ) with equivariant properties (satorras2021n, ; han2022equivariant, ; liu2024segno, ) have become essential tools in modeling complex dynamic systems (karplus2002molecular, ; huang2021equivariant, ; liu2024dynamic, ; lu2024pascl, ) and molecular properties (brandstetter2021geometric, ; meng2024heterogeneous, ) due to their ability to capture symmetrical relationships within data. Equivariant GNNs are particularly effective in scenarios when maintaining consistency under transformations (e.g., rotations, reflections) is crucial. The consistency of equivariant representations allows these models to accurately reflect the inherent properties of the systems they are designed to study, leading to more reliable and insightful predictions, which now is an essential topic in machine learning (battaglia2016interaction, ; thomas2018tensor, ; zhang2022would, ; zhang2023investigating, ; meng2024deep, ).
Despite their successes, existing Equivariant GNNs face notable limitations that constrain their performance and expressiveness. First, the over-smoothing problem (li2018deeper, ; keriven2022not, ; min2020scattering, ; yin2023omg, ). The classical Equivariant GNNs (such as EGNN (satorras2021n, ) and SEGNN (brandstetter2021geometric, )) typically utilize the GNNs (kipf2017semisupervised, ; yin2022generic, ; ai2023gcn, ; shou2023adversarial, ) as a backbone, as well as cooperating coordinates for equivariant representation. However, GNNs usually suffer from the over-smoothing problem, leading to the indistinguishable representation of features and coordinates, which is shown in Fig. 1. Second, gradient explosion or vanishing problems (dasoulas2021lipschitz, ; guo2022orthogonal, ; yinsport, ). Motivated by the notion that larger models are usually more expressive (kaplan2020scaling, ), we tend to explore the expressiveness ability by enlarging the model scale. However, as the network depth increases, gradients can either become excessively large or diminish to near zero, which limits the scalability and depth of GNNs. Third, real-world simulation (norcliffe2020second, ). Most Equivariant GNNs (fuchs2020se, ; thomas2018tensor, ; satorras2021n, ; brandstetter2021geometric, ; liu2023graph, ; liu2022mlp, ) primarily focus on first-order information. However, many physical systems are governed by second-order laws (norcliffe2020second, ). The first-order methods overlook the higher-order interactions that are critical in real-world systems, and limit the representational ability. Even though some works (huang2021equivariant, ; liu2023improving, ) try to introduce second-order information into equivariant learning, they simply focus on the simulation the position in dynamic physical system, ignoring the expressiveness representation for molecular properties prediction.
(a) Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features.
(b) Dirichlet energy $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions.
Figure 1. Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features $\mathbf{H}^{n}$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions $\mathbf{X}^{n}$ propagated through a EGNN (satorras2021n, ), SEGNN (brandstetter2021geometric, ) and GMN (huang2021equivariant, ) on a N-body system dataset, where we give the definition of the Dirichlet energy in Eq. 4.
To address these challenges, we take an insight into the second-order method in Equivariant GNNs and a framework named Dual Second-order Equivariant Graph Ordinary Differential Equation (DuSEGO). The proposed DuSEGO is a universal framework that transfer the first-order Equivariant GNNs to second-order, which aims to enhance Equivariant GNNs expressiveness and address the issues of over-smoothing and gradient stability. First, we prove that the proposed DuSEGO maintains the equivariant properties. Then, we analyze the properties of DuSEGO to reveal how the proposed method solves the over-smoothing and gradient exploding and vanishing problems, facilitating the training of deep GNNs. Extensive experiments on benchmark datasets validate the superiority of the proposed DuSEGO compared to existing baselines.
The main contributions can be summarized as three-fold:
  • The proposed DuSEGO is flexible to accommodate the first-order equivariant graph ODE methods to the second-order, which still maintains the equivariant properties.
  • We analyze the properties of DuSEGO, and demonstrate that the second-order graph ODE can effectively mitigate the over-smoothing, gradient exploding, and vanishing problems.
  • Extensive experiments conducted on various real-world datasets demonstrate the effectiveness of the proposed DuSEGO.
2. Related Work
Equivariant Graph Neural Networks. GNNs (kipf2017semisupervised, ; pang2023sa, ; yin2023messages, ; wang2024degree, ) with equivariant properties have demonstrated their effectiveness in various tasks, especially for 3D molecular graphs and physical symmetry of systems (zhang2023universal, ). The Tensor field networks (TFNs) (thomas2018tensor, ) method was first proposed for 3D point clouds, which only leverage the positional information of nodes without any features. SE(3)-transformers (fuchs2020se, ) deploy attention mechanism into TFNs to improve the model performance. EGNN (satorras2021n, ), a simple but effective form of equivariant graph neural network, was proposed and has been the backbone of the following research. SEGNN (brandstetter2021geometric, ) incorporates geometric and physical properties into the message-passing framework, significantly enhancing the accuracy and predictive capabilities of EGNN in understanding three-dimensional molecular structures. However, these methods focus on first-order velocity information, despite the fact that many physical systems are governed by second-order motion laws. GMN (huang2021equivariant, ) first introduced the second-order information into equivariant learning, and SEGNO (liu2023improving, ) further applied the second-order ODE for solving the physical dynamic system. However, these methods only target the description of positional state information in the physical world and cannot be applied to molecular feature representation learning. The proposed DuSEGO utilizes the second-order ODE to evaluate the evolution of equivariant representation and position, making it suitable for applications in dynamic physical systems and molecular properties prediction.
Graph Ordinary Differential Equations. Graph Ordinary Differential Equations (Graph ODEs) (poli2019graph, ) draw inspiration from Neural Ordinary Differential Equations (chen2018neural, ) and provide a sophisticated mechanism for modeling dynamic changes within graph structures, showing great potential in graph learning (bhattoo2023learning, ; poli2019graph, ; sanchez2019hamiltonian, ; bishnoi2022enhancing, ). A series of works propose to deploy first-order or second-order ODE into GNNs to enhance their expressiveness ability. LG-ODE (huang2020learning, ) learns the dynamics of continuous systems by using models that can handle irregular sampling and incomplete data, improving prediction accuracy and robustness. CG-ODE (huang2021coupled, ) uses coupled ordinary differential equations on graphs to learn and model the dynamics of interacting systems effectively. HOPE (luo2023hope, ) leverages second-order differential equations to capture complex interactions and dynamic behaviors within graph structures. GraphCon (rusch2022graph, ) proposes a novel framework that integrates the dynamics of coupled oscillators with graph theory and analyzes the stability across complex networked systems. However, the above works do not consider the equivariance contained in the graph structure, which is hard to extend into 3D molecular graphs and the physical symmetry of systems.
3. Preliminary
Equivariance Let $T_{g}:X\rightarrow X$ be a set of transformations on $X$ for the abstract group $g\in G$. We say a function $\phi:X\rightarrow Y$ is equivariant to $g$ if there exists an equivalent transformation on its output space $S_{g}:Y\rightarrow Y$ such that: $\phi(T_{g}(\mathbf{x}))=S_{g}(\phi(\mathbf{x}))$. The equivariance properties include translations, rotations, and permutation, which ensures that the predictions of a model are consistent regardless of the orientation or position of the objects within the input data.
  • Translation: Let $\mathbf{x}+g$ be shorthand for $(\mathbf{x_{1}}+g,\cdots,\mathbf{x_{M}}+g)$, then $\phi(x)+g=\phi(x+g)$.
  • Rotation: Let $Q\mathbf{x}$ be shorthand for $(Q\mathbf{x_{1}},\cdots,Q\mathbf{x_{M}})$, then $Q\phi(\mathbf{x})=\phi(Q\mathbf{x})$.
Graph Nerual Networks Given graph $G=(\mathcal{V},\mathcal{E})$, let $\mathbf{h}_{v}^{(k)}$ denote embedding vector for node $v$ at the $l$-th layer. For node $v\in\mathcal{V}$, we aggregate the embeddings from the neighbors of $v$ at layer $l-1$. Then, the node embedding $\mathbf{h}_{v}^{(k)}$ is iteratively updated by combining the embedding of $v$ in the previous layer with the aggregated neighbor embedding. In formulation, $\mathbf{h}_{v}^{l}$ is calculated as follows:
(1) $\mathbf{h}_{v}^{l}=\mathcal{C}^{(k)}\left(\mathbf{h}_{v}^{l-1},\mathcal{A}^{(k)}\left(\left\{\mathbf{h}_{u}^{l-1}\right\}_{u\in\mathcal{N}(v)}\right)\right),$
where $\mathcal{N}(v)$ represents the neighbors of $v$. $\mathcal{A}^{l}$ and $\mathcal{C}^{l}$ represent the aggregation and combination operations at the $l$-th layer, respectively. At last, we summarize all node representations at the $L$-th layer with a readout function into the graph-level representation, which can be formulated as follows:
(2) $\mathbf{z}=F\left(G\right)=\operatorname{READOUT}\left(\left\{\mathbf{h}_{v}^{L}\right\}_{v\in\mathcal{V}}\right),$
where $\mathbf{z}$ is the graph-level representation of $G$ and $\mathbf{\theta_{e}}$ denotes the parameter of our GNN-based encoder. $L$ denotes the number of the graph convolutional layers. The readout function can be implemented in different ways, such as summarizing all nodes’ representations (xu2018powerful, ; yin2022deal, ; yin2023dream, ) or using a virtual node (li2015gated, ; ju2024survey, ; yin2024continuous, ).
Equivariant GNNs Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $v_{i}\in\mathcal{V}$ and edges $e_{ij}\in\mathcal{E}$. In addition to the feature node embeddings $\mathbf{h}_{i}\in\mathbb{R}^{d}$, the $n$-dimensional coordinate $\mathbf{x}_{i}\in\mathbb{R}^{n}$ associated with each of the graph nodes are also considered, where $d$ is the dimension of node feature, and $n$ is the dimension of node position. A GNN with Equivariance $\phi_{\theta}$ with parameter $\theta$ can complete given tasks by message passing layers. Specially, each layer of $\phi_{\theta}$ can be formulated as follows:
(3) $\boldsymbol{m}_{ij}=\mu\left(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},\mathbf{x}_{i}^{l},\mathbf{x}_{j}^{l},e_{ij}\right),\quad\mathbf{x}_{i}^{l+1},\mathbf{h}_{i}^{l+1}=\nu\left(\mathbf{h}_{i}^{l},\mathbf{x}_{i}^{l},\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij}\right),$
where $\mu$ and $\nu$ are the edge message function and node update function, $\mathcal{N}(i)$ are the neighbors of node $i$, $\mathbf{x}_{i}^{l}\in\mathbb{R}^{n}$ and $\mathbf{h}_{i}^{l}\in\mathbb{R}^{d}$ are the updated feature and position of node $i$. For containing E(n) equivariance for each layer, $\mu$ and $\nu$ could be both equivariant (e.g., SEGNN (brandstetter2021geometric, )) or alternatively, $\mu$ is equivariant and $\nu$ is invariant (e.g., EGNN (satorras2021n, )).
Dirichlet Energy As introduced in (rusch2022graph, ), let $\mathbf{X}^{n}$ denote embeddings of node on the $n$-th layer. The Dirichlet energy function is defined as:
(4) $E(\mathbf{X}^{n})\leq C_{1}\exp^{-C_{2}n},$
where $C_{1},C_{2}>0$ are some constant numbers.
4. Method
In this section, we introduce the proposed DuSEGO. In the dynamic systems and molecular properties prediction, we utilize the second-order ODE to simulate the representation and trajectory.
4.1. Second-order Graph ODE
To model high-order correlations in long-term temporal trends,  (rusch2022graph, ) first propose the second-order graph ODE, which is represented as:
(5) $\mathbf{X}^{{}^{\prime\prime}}=\sigma(\mathbf{F}_{\theta}(\mathbf{X},t))-\gamma\mathbf{X}-\alpha\mathbf{X}^{{}^{\prime}},$
where $\left(\mathbf{F}_{\theta}(\mathbf{X},t)\right)_{i}=\mathbf{F}_{\theta}\left(\mathbf{X}_{i}(t),\mathbf{X}_{j}(t),t\right)$ is a learnable coupling function with parameters $\theta$. $\sigma$ is the mapping function, which is implemented with multiplayer perceptions. Due to the unavailability of an analytical solution for Eq. 5, GraphCON (rusch2022graph, ) addresses it through an iterative numerical solver employing a suitable time discretization method. GraphCON utilizes the IMEX (implicit-explicit) time-stepping scheme, an extension of the symplectic Euler method (book) that accommodates systems with an additional damping term.
(6) $\mathbf{Y}^{n}=\mathbf{Y}^{n-1}+\Delta t$ $\left[\sigma(\mathbf{F}_{\theta}(\mathbf{X}^{n-1},t^{n-1}))-\gamma\mathbf{X}^{n-1}-\alpha\mathbf{Y}^{n-1}\right],$
$\mathbf{X}^{n}=$ $\mathbf{X}^{n-1}+\Delta t\mathbf{Y}^{n},\,n=1,\cdots,N,$
where $\Delta t>0$ is a fixed time-step and $\mathbf{Y}^{n}$, $\mathbf{X}^{n}$ denote the hidden node features at time $t^{n}=n\Delta t$.
4.2. DuSEGO
To efficiently evaluate the graph representation and the coordinate information, we present the framework in a graph dynamical system by following the second-order ODEs:
(7) $\mathbf{X}^{\prime\prime}=\sigma(\mathbf{F}_{\theta}(\mathbf{X},\mathbf{H},t))-\gamma_{1}\mathbf{X}-\alpha\mathbf{X}^{\prime},\quad\mathbf{H}^{\prime\prime}=\sigma(\mathbf{F}_{\theta}(\mathbf{X},\mathbf{H},t))-\gamma_{2}\mathbf{H}-\alpha\mathbf{H}^{\prime},$
where $\mathbf{X}$ is the coordinate embeddings and $\mathbf{H}$ is the node features. Similar with Eq. 6, we have:
(8) $\mathbf{Y}^{n}=\mathbf{Y}^{n-1}+\Delta t\left[\sigma(\mathbf{F}_{\theta}(\mathbf{X}^{n-1},\mathbf{H}^{n-1},t^{n-1}))-\gamma_{1}\mathbf{X}^{n-1}-\alpha\mathbf{Y}^{n-1}\right],\mathbf{X}^{n}=\mathbf{X}^{n-1}+\Delta t\mathbf{Y}^{n},\,n=1,\cdots,N,$
and
(9) $\mathbf{U}^{n}=\mathbf{U}^{n-1}+\Delta t\left[\sigma(\mathbf{F}_{\theta}(\mathbf{X}^{n-1},\mathbf{H}^{n-1},t^{n-1}))-\gamma_{2}\mathbf{H}^{n-1}-\alpha\mathbf{U}^{n-1}\right],\mathbf{H}^{n}=\mathbf{H}^{n-1}+\Delta t\mathbf{U}^{n},\,n=1,\cdots,N,$
To keep the Equivariant properties of DuSEGO, the proposed method allows for any Equivariant GNNs to be used as $\mathbf{F}_{\theta}$.
Choice of function $\mathbf{F}_{\theta}$. Our framework allows for any learnable equivariant GNNs to be used as $\mathbf{F}_{\theta}$. In this paper, we focus on the two particularly popular choices:
Choice 1: Equivariant Graph Neural Networks (EGNNs):
(10) $\mathbf{F}_{\theta}(\mathbf{X}^{l},\mathbf{H}^{l},t^{l})=\textbf{h}_{i}^{l+1},\;\mathbf{x}_{i}^{l+1}=\phi_{h}(\textbf{h}_{i}^{l},\textbf{m}_{i}),\;\mathbf{x}_{i}^{l}+C\sum_{j\neq i}(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l})\phi_{x}(\mathbf{m}_{ij}),$
where $\mathbf{m}_{ij}=\phi_{e}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},||\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}||^{2},a_{ij})$ and $\textbf{m}_{i}=\sum_{j\neq i}\textbf{m}_{ij}$. $\phi_{e}$ and $\phi_{h}$ are the edge and node operations respectively which are commonly approximated by Multilayer Perceptrons (MLPs). Nodes $j$ are the neighbors of node $i$.
Choice 2: Steerable E(3) Equivariant Graph Neural Networks (SEGNN):
(11) $\mathbf{F}_{\theta}(\mathbf{X}^{l},\mathbf{H}^{l},t^{l})=\phi_{h}(\mathbf{h}_{i}^{l},\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij},\mathbf{a}_{i}),\;\mathbf{m}_{ij}=\phi_{m}(\mathbf{h}_{i},\mathbf{h}_{j},||\mathbf{x}_{j}-\mathbf{x}_{i}||^{2},\mathbf{a}_{ij}),$
where $||\mathbf{x}_{j}-\mathbf{x}_{i}||^{2}$ is the squared relative distance between two nodes $v_{i}$ and $v_{j}$, $\phi_{m}$ and $\phi_{h}$ are O(3) steerable MLPs, $\mathbf{a}_{ij}$ and $\mathbf{a}_{i}$ are steerable edge and node attributes, $\mathbf{a}_{i}=\frac{1}{|\mathcal{N}(I)|}\sum_{j\in\mathcal{N}(i)}\mathbf{a}_{ij}$. $\mathcal{N}(i)$ denotes the neighbors of node $v_{i}$.
Steady States of DuSEGO. It is straightforward to see that the steady states $\mathbf{X}^{*}$,$\mathbf{H}^{*}$ of the DuSEGO dynamical system with an autonomous function $\mathbf{F}_{\theta}=\mathbf{F}_{\theta}(\mathbf{X})$ are given by $\mathbf{Y}^{*}\equiv 0$ and $\mathbf{U}^{*}\equiv 0$.
(12) $\mathbf{\mathbf{X}^{*}}=\frac{\Delta t}{\gamma_{1}}\sigma(\mathbf{F}_{\theta}(\mathbf{X}^{*})),\;\mathbf{\mathbf{H}^{*}}=\frac{\Delta t}{\gamma_{2}}\sigma(\mathbf{F}_{\theta}(\mathbf{H}^{*})).$
Thus, this interpretation of DuSEGO clearly brings out its relationship with equivariant GNNs. Unlike in standard multi-layer GNNs of the generic that can be thought of as steady states of the underlying ODEs, DuSEGO evolves the underlying node features dynamically in time. Interpreting the multiple GNN layers as iterations at times $t_{n}=n\Delta t$, we observe that the node features in DuSEGO follow the trajectories of the corresponding dynamical system and can explore a richer sampling of the underlying latent feature space, leading to possibly greater expressive power than equivariant GNNs, which might remain in the vicinity of steady states.
5. DuSEGO Properties Analysis
To gain some insights into the functioning of DuSEGO, we start by setting the hyperparameter $\gamma=1$ and assuming that the 1-neighborhood coupling $F_{\theta}$ is given by either the GAT or GCN type coupling functions. In this case, the underlying ODEs take the following node-wise form,
Proposition 0.
Suppose the backbone Equivariant GNN $F_{\theta}$ of DuSEGO is E(3) equivariant and translation- invariant, the output $\mathbf{H}$ and $\mathbf{X}$ is E(3)-equivariant.
Proof: Due to $F_{\theta}$ is E(3) equivariant, then we have:
(13) $F_{\theta}(\mathbf{X}+g,\mathbf{H}+g)=F_{\theta}(\mathbf{X},\mathbf{H})+g,\quad F_{\theta}(Q\mathbf{X},Q\mathbf{H})=QF_{\theta}(\mathbf{X},\mathbf{H}).$
For the proposed DuSEGO, we ignore the $\sigma$ and set $\gamma_{1}=1$, $\gamma_{2}=1$ for simplification,
$F_{\theta}(\mathbf{X}+g,\mathbf{H}+g)-(\mathbf{X}+g)-\alpha((\mathbf{X}+g)^{\prime})=F_{\theta}(\mathbf{X},\mathbf{H})+g-\mathbf{X}-g-\alpha\mathbf{X}^{\prime}=F_{\theta}(\mathbf{X},\mathbf{H})-\mathbf{X}-\alpha\mathbf{X}^{\prime}=\mathbf{X}^{\prime\prime},$
$F_{\theta}(\mathbf{X}+g,\mathbf{H}+g)-(\mathbf{H}+g)-\alpha((\mathbf{H}+g)^{\prime})=F_{\theta}(\mathbf{X},\mathbf{H})+g-\mathbf{H}-g-\alpha\mathbf{H}^{\prime}=F_{\theta}(\mathbf{X},\mathbf{H})-\mathbf{H}-\alpha\mathbf{H}^{\prime}=\mathbf{H}^{\prime\prime}.$
Similarly, we have:
(14) $F_{\theta}(Q\mathbf{X},Q\mathbf{H})-Q\mathbf{X}-\alpha Q\mathbf{X}^{\prime}=QF_{\theta}(\mathbf{X},\mathbf{H})-Q\mathbf{X}-\alpha Q\mathbf{X}^{\prime}=Q\mathbf{X}^{\prime\prime},$
(15) $F_{\theta}(Q\mathbf{X},Q\mathbf{H})-Q\mathbf{H}-\alpha Q\mathbf{H}^{\prime}=QF_{\theta}(\mathbf{X},\mathbf{H})-Q\mathbf{H}-\alpha Q\mathbf{H}^{\prime}=Q\mathbf{H}^{\prime\prime}.$
Therefore, the output $\mathbf{H}$ and $\mathbf{X}$ also follows E(3) equivariant.
Proposition 0.
Let $\mathbf{X}_{n}$, $\mathbf{Y}_{n}$ be the node features, and $\mathbf{H}_{n}$, $\mathbf{U}_{n}$ be the node coordinate, generated by DuSEGO. We assume that $\Delta t\ll 1$ is chosen to be sufficiently small. Then, the gradient of the loss function $\mathcal{L}$ with respect to any learnable weight parameter $\mathbf{w}_{k}^{l}$, for some $1\leq k\leq v$ and $1\leq l\leq N$ is bounded as:
(16) $\left|\frac{\partial\mathcal{L}}{\partial\mathbf{w}_{k}^{l}}\right|\leq$ $\frac{\hat{D}\Delta t}{v}\Bigg{(}\max\limits_{x}|\sigma^{\prime}(x)|\cdot(1+\Gamma_{x}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\mathbf{X}_{i}^{0}|+|\mathbf{Y}_{i}^{0}|\Big{)}$
$+\max\limits_{h}|\sigma^{\prime}(h)|\cdot(1+\Gamma_{h}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\mathbf{H}_{i}^{0}|+|\mathbf{U}_{i}^{0}|\Big{)}\Bigg{)}$
$+\frac{\hat{D}\Delta t}{v}\Bigg{(}\max\limits_{x}|\sigma^{\prime}(x)|\cdot(1+\Gamma_{x}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\bar{\mathbf{X}}_{i}|+\max\limits_{x}|\sigma(x)|\sqrt{N\Delta t}\Big{)}^{2}$
$+\max\limits_{h}|\sigma^{\prime}(h)|\cdot(1+\Gamma_{h}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\bar{\mathbf{H}}_{i}|+\max\limits_{h}|\sigma(h)|\sqrt{N\Delta t}\Big{)}^{2}\Bigg{)}.$
where $\hat{D}=\max\limits_{i,j\in\mathcal{V}}\frac{1}{\sqrt{d_{i}d_{j}}}$, and $\Gamma_{x}:=6+4\max\limits_{x}|\sigma^{{}^{\prime}}(x)|\cdot\hat{D}\max\limits_{1\leq n\leq T}||\mathbf{w}^{n}||_{1}$ and $\Gamma_{h}:=6+4\max\limits_{h}|\sigma^{{}^{\prime}}(h)|\cdot\hat{D}\max\limits_{1\leq n\leq T}||\mathbf{w}^{n}||_{1}$. $d_{i}$ is the degree of node $i$, $\mathbf{\bar{X}}_{i}$ is the label of node $i$ and $\mathbf{\bar{H}}_{i}$ is the coordinate of node $i$.
Proof: From (rusch2022graph, ), we have:
(17) $\left\|\frac{\partial\mathbf{H^{N}}}{\partial\mathbf{H}^{l}}\right\|_{\infty}\leq 1+(N-l)\Gamma_{h}\Delta t\leq 1+N\Gamma_{h}\Delta t,\quad\left\|\frac{\partial\mathbf{X^{N}}}{\partial\mathbf{X}^{l}}\right\|_{\infty}\leq 1+(N-l)\Gamma_{x}\Delta t\leq 1+N\Gamma_{x}\Delta t,$
where $\Gamma_{h}=6+4\max\limits_{h}|\sigma^{\prime}(h)|\hat{D}\max\limits_{1\leq n\leq T}\|\mathbf{w}^{n}\|_{1}$, $\Gamma_{x}=6+4\max\limits_{x}|\sigma^{\prime}(x)|\hat{D}\max\limits_{1\leq n\leq T}\|\mathbf{w}^{n}\|_{1}$, and $\hat{D}=\max\limits_{i,j\in\mathcal{V}}\frac{1}{\sqrt{d_{i}d_{j}}}$. By the chain rule, the gradient of $\mathbf{w}_{k}^{l}$ is:
(18) $\frac{\partial\mathcal{L}}{\partial\mathbf{w}_{k}^{l}}=\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{N}}\frac{\partial\mathbf{H}^{N}}{\partial\mathbf{H}^{l}}\frac{\partial\mathbf{H}^{l}}{\partial\mathbf{w}_{k}^{l}}+\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{N}}\frac{\partial\mathbf{X}^{N}}{\partial\mathbf{X}^{l}}\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{w}_{k}^{l}},$
where $\frac{\partial\mathbf{H}^{N}}{\partial\mathbf{H}^{l}}=\Pi^{N}_{n=l+1}\frac{\partial\mathbf{H}^{n}}{\partial\mathbf{H}^{n-1}}$, $\frac{\partial\mathbf{X}^{N}}{\partial\mathbf{X}^{l}}=\Pi^{N}_{n=l+1}\frac{\partial\mathbf{X}^{n}}{\partial\mathbf{X}^{n-1}}$. With Eq. 51 in (rusch2022graph, ), we directly have:
(19) $\left\|\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{N}}\right\|_{\infty}\leq\frac{1}{v}\left(\max\limits_{1\leq i\leq v}(|\mathbf{X}_{i}^{0}|+|\mathbf{Y}_{i}^{0}|)+\max\limits_{1\leq i\leq v}|\bar{\mathbf{X}}_{i}|+\max\limits_{x}|\sigma(x)|\sqrt{N\Delta t}\right),$
(20) $\left\|\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{N}}\right\|_{\infty}\leq\frac{1}{v}\left(\max\limits_{1\leq i\leq v}(|\mathbf{H}_{i}^{0}|+|\mathbf{U}_{i}^{0}|)+\max\limits_{1\leq i\leq v}|\bar{\mathbf{H}}_{i}|+\max\limits_{h}|\sigma(h)|\sqrt{N\Delta t}\right),$
where $\bar{\mathbf{X}}$ and $\bar{\mathbf{H}}$ are the ground truth vector. Similarly, we can readily calculate that:
(21) $\left\|\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{w}_{k}^{l}}\right\|_{\infty}\leq\Delta t\max\limits_{x}|\sigma^{\prime}(x)|\hat{D}\left(\max\limits_{1\leq i\leq v}(|\mathbf{X}_{i}^{0}|+|\mathbf{Y}_{i}^{0}|)+\max\limits_{1\leq i\leq v}|\bar{\mathbf{X}}_{i}|+\max\limits_{x}|\sigma(x)|\sqrt{N\Delta t}\right),$
(22) $\left\|\frac{\partial\mathbf{H}^{l}}{\partial\mathbf{w}_{k}^{l}}\right\|_{\infty}\leq\Delta t\max\limits_{h}|\sigma^{\prime}(h)|\hat{D}\left(\max\limits_{1\leq i\leq v}(|\mathbf{H}_{i}^{0}|+|\mathbf{U}_{i}^{0}|)+\max\limits_{1\leq i\leq v}|\bar{\mathbf{H}}_{i}|+\max\limits_{h}|\sigma(h)|\sqrt{N\Delta t}\right).$
Therefore, by multiplying Eq. 17,  19-22, we have:
(23) $\left|\frac{\partial\mathcal{L}}{\partial\mathbf{w}_{k}^{l}}\right|\leq$ $\frac{\hat{D}\Delta t}{v}\Bigg{(}\max\limits_{x}|\sigma^{\prime}(x)|\cdot(1+\Gamma_{x}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\mathbf{X}_{i}^{0}|+|\mathbf{Y}_{i}^{0}|\Big{)}$
$+\max\limits_{h}|\sigma^{\prime}(h)|\cdot(1+\Gamma_{h}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\mathbf{H}_{i}^{0}|+|\mathbf{U}_{i}^{0}|\Big{)}\Bigg{)}$
$+\frac{\hat{D}\Delta t}{v}\Bigg{(}\max\limits_{x}|\sigma^{\prime}(x)|\cdot(1+\Gamma_{x}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\bar{\mathbf{X}}_{i}|+\max\limits_{x}|\sigma(x)|\sqrt{N\Delta t}\Big{)}^{2}$
$+\max\limits_{h}|\sigma^{\prime}(h)|\cdot(1+\Gamma_{h}N\Delta t)\Big{(}\max\limits_{1\leq i\leq v}|\bar{\mathbf{H}}_{i}|+\max\limits_{h}|\sigma(h)|\sqrt{N\Delta t}\Big{)}^{2}\Bigg{)}.$
6. Experiments
To verify the effectiveness of DuSEGO in enhancing the performance of Equivariant GNN, we follow the experimental settings in (satorras2021n, ) and evaluate it on three different tasks, i.e., modeling a dynamical system, learning unsupervised representations of graphs in a continuous latent space, and predicting molecular chemical properties.
6.1. Modelling a Dynamic System
In this section, we evaluate DuSEGO on one simulated N-body system named charged particles, where we aim to forecast the position of particles across different time steps (satorras2021n, ).
6.1.1. Dataset
Building on the work of (fuchs2020se, ), (satorras2021n, ) extended the Charged Particles N-body experiment introduced in (kipf2018neural, ), where the system consists of five particles, each defined by charge, position, and velocity in three-dimensional space. In this setup, particles in each graph are fully connected, with the initial velocity norm included as an additional node feature. Using the N-body system code from (satorras2021n, ), we generated a new dataset for the charged N-body system across 1,000, 1,500, and 2,000 timesteps. The initial positions of particles were sampled from a Gaussian distribution with $\mu=0$ and $\sigma=0.5$, and the initial velocity was represented as a random vector with a norm of 0.5. The dataset is split into training, validation, and testing sets, containing 10,000, 2,000, and 2,000 samples, respectively.
Table 1. Mean squared error (×$\left(\times 10^{-2}\right)$ for the future position estimation under different timesteps in the N-body system experiment, and forward time in seconds for a batch size of 100 samples running on NVIDIA A100 GPUs. Results averaged across 5 runs. We report both mean and standard deviation of MSE in the table. The best performance is bold.
Methods MSE Forward Time/ms
1000 ts 1500 ts 2000 ts
Linear 6.831 $\pm$ 0.016 20.012 $\pm$ 0.029 39.513 $\pm$ 0.061 0.1
SE(3) Trans 2.483 $\pm$ 0.099 18.891 $\pm$ 0.287 36.730 $\pm$ 0.381 114.6
TFN 1.544 $\pm$ 0.231 11.116 $\pm$ 2.825 23.823 $\pm$ 3.048 27.2
GNN 1.077 $\pm$ 0.004 5.059 $\pm$ 0.250 10.591 $\pm$ 0.352 2.1
Radial Field 1.060 $\pm$ 0.007 12.514 $\pm$ 0.089 26.388 $\pm$ 0.331 3.0
EGNN 0.716 $\pm$ 0.029 2.201 $\pm$ 0.081 4.049 $\pm$ 0.103 5.5
SEGNN 0.481 $\pm$ 0.016 1.552 $\pm$ 0.061 3.294 $\pm$ 0.095 24.5
GraphCon 1.015 $\pm$ 0.011 3.148 $\pm$ 0.091 6.133 $\pm$ 0.143 10.8
SEGNO 0.976 $\pm$ 0.012 2.994 $\pm$ 0.059 7.453 $\pm$ 0.181 21.2
GMN 0.701 $\pm$ 0.018 1.956 $\pm$ 0.096 3.939 $\pm$ 0.151 6.2
DuSEGO-EGNN 0.674 $\pm$ 0.014 1.912 $\pm$ 0.055 3.876 $\pm$ 0.079 13.9
DuSEGO-SEGNN 0.449 $\pm$ 0.012 1.427 $\pm$ 0.047 3.122 $\pm$ 0.081 45.7
6.1.2. Baselines.
Two types of baseline methods are deployed for comparison: (1) Methods that incorporate first-order information, including the Linear model, Graph Neural Networks (GNN), SE(3) Transformer (fuchs2020se, ), Tensor Field Networks (TFN) (thomas2018tensor, ), Radial Field (RF) (kohler2019equivariant, ), Equivariant Graph Neural Networks (EGNN) (satorras2021n, ), and Steerable E(3) Equivariant Graph Neural Networks (SEGNN) (brandstetter2021geometric, ); (2) Models that incorporate second-order information, including GraphCon (rusch2022graph, ), Graph Mechanics Network (GMN) (huang2021equivariant, ), and Second-order Equivariant Graph Neural Ordinary Differential Equation (SEGNO) (liu2023improving, ).
6.1.3. Experimental Setting.
We build on the experimental framework introduced in (liu2023improving, ), aiming to estimate the positions of all particles after 1,000, 1,500, and 2,000 timesteps. We re-evaluate all baseline methods under identical experimental settings on NVIDIA A100 GPUs, gathering results across five runs to ensure fair comparisons. The experimental setup is as follows: Adam optimizer with a learning rate of $5\times 10^{-4}$, batch size of 100, 10,000 epochs, hidden embedding dimension of 64, weight decay of $1\times 10^{-12}$, and four GNN layers. Specifically, for the GMN method, we set the numbers of isolated particles, sticks, and hinges to $(5,0,0)$. For the SE(3)-transformer and TFN, we set the representation degrees to 3 and 2, respectively. An early stopping criterion is applied, halting training if no improvement in validation loss is observed within 50 epochs.
6.1.4. Implementation details.
We employ EGNN with the extension that includes velocity and SEGNN as the backbone of DuSEGO, named DuSEGO-EGNN and DuSEGO-SEGNN. Especially, EGNN and SEGNN are implemented based on the implementation details provided in (satorras2021n, ; brandstetter2021geometric, ), which are used as the coupling function for message passing (rusch2022graph, ). Then, we use the ODE solvers provided in torchdiffeq  [^1] https://github.com/rtqichen/torchdiffeq library for solving Eq 8 and Eq. 9. Additionally, we provide the forward time in seconds for each method for a batch of 100 samples.
(a) GMN
(b) EGNN
(c) SEGNN
(d) SEGNO
(e) DuSEGO-EGNN
(f) DuSEGO-SEGNN
Figure 2. Visualization of ground truths (Blue) and prediction results (Red) for EGNN, SEGNN, GMN, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN under 1000 timesteps on the N-body system dataset.
Figure 3. Data efficiency comparisons among EGNN, SEGNN, GMN, GraphCon, SEGNO, DuSEGO-EGNN , and DuSEGO-SEGNN on the N-body dataset.
(a) Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features.
(b) Dirichlet energy $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions.
Figure 4. Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features $\mathbf{H}^{n}$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions $\mathbf{X}^{n}$ propagated through a EGNN and DuSEGO-EGNN for two different values of $\alpha=0,1$ and $\gamma_{1}=\gamma_{2}=1$ in Eq. 8 and Eq. 9.
6.1.5. Results.
The results for all baseline methods, DuSEGO-EGNN, and DuSEGO-SEGNN on the N-body system dataset are shown in Table 1. We observe that DuSEGO significantly enhances the performance of EGNN and SEGNN, with DuSEGO-SEGNN outperforming all other baseline methods and DuSEGO-EGNN under the same experimental conditions. Specifically, DuSEGO-EGNN and DuSEGO-SEGNN reduce the mean squared error (MSE) of SEGNN and EGNN by an average of 7.73% and 6.65%, respectively, across the three timesteps, indicating substantial improvement. Notably, the relative improvement for DuSEGO-EGNN increases from 5.86% at 1,000 timesteps to 13.13% at 1,500 timesteps, highlighting the strong generalization ability of DuSEGO. Additionally, Figure 2 presents a visualization of GMN, EGNN, SEGNN, GMN, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN in estimating particle positions for the N-body system experiment at 1,000 timesteps, with MSE values between predictions and ground truths reported. This visualization further shows that DuSEGO-SEGNN and DuSEGO-EGNN outperform EGNN and SEGNN, respectively, aligning with the performance results in Table 1.
It is crucial to verify that the physical principles identified from a limited number of experiments are sufficiently general to explain phenomena on a broader scale. To this end, we assess the performance of EGNN, SEGNN, GMN, GraphCon, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN across various training set sizes, ranging from 500 to 3,000 samples in the N-body system dataset. In this evaluation, the timesteps of particle positions are set to 1,000. As shown in Fig. 3, DuSEGO-EGNN and DuSEGO-SEGNN, similar to the GMN method, demonstrate more stable performance than other baseline methods, emphasizing their data efficiency. This result suggests that data efficiency, a notable strength of DuSEGO, is essential for effectively learning to model physical systems.
To further investigate the over-smoothing issue in EGNN, we implement EGNN and DuSEGO-EGNN with two different values of $\alpha=0,1$ and set $\gamma_{1}=\gamma_{2}=1$ in Eq. 8 and Eq. 9 on the N-body system dataset at 1,000 timesteps. As shown in Fig. 4, the over-smoothing problem in EGNN becomes apparent as the number of EGNN layers increases. The differences between node features, such as $\mathbf{h}_{i}^{n}$ and $\mathbf{h}_{j}^{n}$, as well as node positions, like $\mathbf{x}_{i}^{n}$ and $\mathbf{x}_{j}^{n}$, gradually decrease, indicating that node features and positions are converging toward a constant value. However, with the integration of DuSEGO into EGNN, the over-smoothing issue is effectively mitigated, as $\mathbf{E}\left(\mathbf{H}^{n}\right)$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ remain stable within a reasonable range, even as the number of EGNN layers increases.
Additionally, we compare the performance of EGNN and DuSEGO-EGNN as the number of EGNN layers increases (up to 16). As shown in Table 2, the MSE for EGNN decreases from layer 4 to layer 8. However, at 12 layers, the gradient vanishing issue emerges, preventing EGNN from producing accurate results. This confirms that gradient vanishing becomes a limitation as the depth of EGNN increases. In contrast, the MSE for DuSEGO-EGNN continues to decrease with additional layers across different timesteps, and DuSEGO-EGNN maintains stable performance, supporting our hypothesis that integrating EGNN within DuSEGO helps mitigate the gradient vanishing problem.
Table 2. Mean squared error (×$\left(\times 10^{-3}\right)$ for the future position estimation under different timesteps in the N-body system experiment between EGNN and DuSEGO-EGNN under different layers. Results averaged across 5 runs.
Methods MSE
1000 ts 1500 ts 2000 ts
EGNN-layers=4 7.16 $\pm$ 0.29 22.01 $\pm$ 0.81 40.49 $\pm$ 1.03
EGNN-layers=8 6.82 $\pm$ 0.23 21.39 $\pm$ 1.46 38.97 $\pm$ 2.77
EGNN-layers=12 NAN NAN NAN
EGNN-layers=16 NAN NAN NAN
DuSEGO-EGNN -layers=4 6.74 $\pm$ 0.14 19.12 $\pm$ 0.55 38.76 $\pm$ 0.79
DuSEGO-EGNN -layers=8 6.68 $\pm$ 0.12 18.42 $\pm$ 0.78 37.46 $\pm$ 1.13
DuSEGO-EGNN -layers=12 6.62 $\pm$ 0.11 17.62 $\pm$ 0.68 36.57 $\pm$ 1.62
DuSEGO-EGNN -layers=16 6.49 $\pm$ 0.18 17.03 $\pm$ 0.99 35.91 $\pm$ 2.01
6.2. Graph Autoencoder
Subsequently, we compare DuSEGO with the state-of-art baseline methods for the Graph Autoencoder task, where we aim to estimate the adjacency matrix ${A}_{ij}$ of a graph $\mathcal{G}$ generated from the graph autoencoder and minimize the binary cross entropy between the estimated adjacency matrix $\hat{A}_{ij}$ and the ground truth $A_{ij}$.
6.2.1. Dataset.
Two widely used datasets are deployed in this section: (1) Community  (kipf2016variational, ) dataset, generated using the Erdos-Renyi model (erdds1959random, ), consists of random graphs where nodes are connected with a predefined probability, capturing the diversity in the density and structure of various communities. This model, denoted as $G(n,p)$, generates networks by independently linking pairs of nodes with probability $p$; (2) Erdos & Renyi (janson2011random, ) dataset, based on the foundational random graph generation model (janson2011random, ; you2018graphrnn, ) and also represented as $G(n,p)$, includes a collection of graphs that exhibit the randomness and variability of network structures.
Following (satorras2021n, ), we generate the Community and Erdos & Renyi datasets (kohler2019equivariant, ; kipf2018neural, ) using the code provided in (kipf2018neural, ). Specially, we sample 50,000 graphs for training, 5,000 for validation, and 5,000 for testing, where we set $p=0.1$ and $6\leq n\leq 10$ for Erdos&Renyi, and $p=0.15$ and $6\leq n\leq 11$ for Community, respectively. Additionally, 20 communities are divided in the Community dataset, and each graph is defined as an adjacency matrix $A\in\{0,1\}^{M\times M}$ in these two datasets.
Table 3. % Error and F1 scores for the test partition on the Graph Autoencoding experiment in the Community and Erdos & Renyi datasets. Results averaged across 5 runs. The best performance is bold.
Methods Community Erdos & Renyi
F1 % Error F1 % Error
GNN 0.957 3.86 0.776 6.44
GNN-Noise 0.971 2.54 0.919 1.79
RFN 0.973 1.81 0.921 1.45
EGNN 0.981 1.58 0.933 1.26
GraphCon 0.960 3.79 0.873 3.77
SEGNO 0.980 1.66 0.920 1.43
GMN 0.979 1.74 0.918 1.46
DuSEGO-EGNN 0.984 1.48 0.936 1.18
Table 4. % Error and F1 scores for the test partition on the Graph Autoencoder experiment in the Erdos&Renyi and Community datasets between EGNN and DuSEGO-EGNN under different layers. Results averaged across 5 runs.
Methods Community Erdos & Renyi
F1 % Error F1 % Error
EGNN-layers=4 0.981 1.58 0.933 1.26
EGNN-layers=6 0.977 2.04 0.929 1.39
EGNN-layers=8 Inf Inf 0.925 1.46
EGNN-layers=12 Inf Inf Inf Inf
EGNN-layers=16 Inf Inf Inf Inf
DuSEGO-EGNN -layers=4 0.984 1.48 0.936 1.18
DuSEGO-EGNN -layers=6 0.980 1.76 0.957 0.77
DuSEGO-EGNN -layers=8 0.981 1.62 0.963 0.63
DuSEGO-EGNN -layers=12 0.982 1.57 0.959 0.79
DuSEGO-EGNN -layers=16 0.982 1.58 0.966 0.59
6.2.2. Baselines
Two types of baseline methods are deployed for comparison: (1) Methods that contains first-order information including GNN, GNN-Noise, RF (kohler2019equivariant, ), EGNN (satorras2021n, ); (2) Methods that contains second-order information including GraphCon (rusch2022graph, ), GMN (huang2021equivariant, ) and SEGNO (liu2023improving, ).
6.2.3. Experimental Setting.
The noise generated from $\mathcal{N}(\mathbf{0},\sigma\mathbf{I})\in\mathbb{R}^{M\times n}$ will be used as input for all baseline methods. The output will then be employed to embed the graph and fed into the graph decoder to generate an estimated adjacency matrix, $\hat{A}_{ij}$. We rerun all baseline methods under the same experimental settings on NVIDIA A100 GPUs and collect the results across five runs for a fair comparison. The experimental settings are as follows: Adam optimizer with learning rate ${1\times 10^{-3}}$, the number of epochs 1,000, batch size 256, hidden embedding dim 64, weight decay ${1\times 10^{-12}}$, layer number 4, and dimensions for the embedding space 8. Velocity is not considered in the Graph Autoencoder task, which is set to None in the EGNN, GMN, SEGNO and DuSEGO-EGNN.
6.2.4. Implementation details.
We build upon the experimental setting introduced in (satorras2021n, ). and employ EGNN as the backbone of DuSEGO, named DuSEGO-EGNN. F1 score of the edge classification and % Error, defined as the percentage of incorrectly predicted edges relative to the total number of potential edges, are reported here. EGNN is implemented based on the implementation details provided in (satorras2021n, ), which is used as the coupling function for message passing (rusch2022graph, ). Then, we use the ODE solvers provided in torchdiffeq library for solving Eq 8 and Eq. 9.
6.2.5. Results.
Table 3 depicts the overall results of all methods on the Community and Erdos & Renyi datasets. It is obvious that DuSEGO-EGNN can outperform all the baseline methods across these two datasets, supporting the general effectiveness of DuSEGO. It is also worth noting that DuSEGO-EGNN utilizes EGNN as its backbone, the second-order information in DuSEGO-EGNN helps EGNN diminish the % Error and increase the F1 score on these two datasets. Additionally, the experimental results demonstrate that DuSEGO-EGNN remains E(n)-equivariant to the noise distribution, provides the best reconstruction of the adjacency matrix ${A}_{ij}$ on a large-scale graph dataset, and enhances performance when embedding graphs into a continuous space as a set of nodes in dimension n.
Then, we investigate the performance of DuSEGO-EGNN as the number of layers increases (up to 16) compared to EGNN, as shown in Table 4. We observe that increasing depth worsens the F1 score and % Error for EGNN. Additionally, when the number of layers reaches 8 on the Community dataset and 12 on the Erdos & Renyi dataset, gradient explosion occurs, causing EGNN to produce infinite values. These results confirm that gradient explosion happens as EGNN depth increases. However, in contrast to EGNN, DuSEGO-EGNN remains stable on both the Erdos & Renyi and Community datasets, even with up to 16 layers, supporting our hypothesis that gradient explosion is mitigated when EGNN is implemented within DuSEGO.
6.3. Molecular Data
Finally, we evaluate DuSEGO with the state-of-art baseline methods in molecular data, where we aim to predict molecular chemical properties on 3D molecular graphs.
6.3.1. Dataset.
The QM9 dataset is widely used in computational chemistry, which is specifically designed for the development and testing of quantum chemistry methods. It consists of approximately 134,000 stable small organic molecules made up of C, H, O, N, and F. Each molecule in the dataset contains up to nine non-hydrogen atoms. The dataset includes a diverse set of molecular properties that have been pre-computed using high-level quantum mechanical theories. We imported the QM9 dataset partitions from (anderson2019cormorant, ) and sample 100K molecules for training, 18K for validation and 13K for testing (brandstetter2021geometric, ). Additionally all properties have been normalized by substracting the mean and dividing by the Mean Absolute Deviation (satorras2021n, ).
6.3.2. Baselines.
Various widely used methods for molecule property prediction are deployed for comparison, such as DimeNet++ (gasteiger_dimenetpp_2020, ), SphereNet (liu2021spherical, ), PaiNN (schutt2021equivariant, ), PamNet (zhang2023universal, ) and LEFTNet (du2024new, ) etc.
Table 5. Performance comparison on the QM9 dataset. Numbers are reported for Mean Absolute Error (MAE) between model predictions and ground truth. ‘-’ denotes no results are reported in the referred papers for the corresponding properties. The best performance is bold.
Task $\alpha$ $\Delta\epsilon$ $\epsilon_{\text{HOMO}}$ $\epsilon_{\text{LUMO}}$ $\mu$ $C_{v}$ $G$ $H$ $R^{2}$ $U$ $U_{0}$ ZPVE
Units bohr3 meV meV meV D cal/mol K meV meV bohr3 meV meV meV
NMP .092 69 43 38 .030 .040 19 17 .180 20 20 1.50
SchNet .297 71 49 41 .049 .043 19 21 .127 24 19 2.03
Cormorant .085 61 34 38 .038 .026 20 21 .961 21 22 2.02
L1Net .088 68 46 35 .043 .031 14 14 .354 21 13 1.56
LieConv .084 49 30 25 .032 .038 22 24 .800 19 19 2.28
TFN .223 58 40 38 .064 .101 - - - - - -
SE(3)-Tr. .142 53 35 33 .051 .054 - - - - - -
DimeNet++ .047 45 27 24 .028 .027 11 11 .573 11 11 1.49
PaiNN .049 49 29 23 .029 .028 11 10 .088 10 10 1.46
SphereNet .056 54 33 24 .031 .027 10 10 .472 10 9 1.44
PAMNet .059 48 28 23 .027 .025 12 12 .990 12 12 1.47
LEFTNet .046 44 27 22 .022 .020 14 14 .169 15 14 1.61
EGNN .071 48 29 25 .028 .032 12 12 .103 12 12 1.60
SEGNN .077 46 28 24 .033 .035 12 11 .899 11 12 1.65
SEGNO .243 86 60 52 .173 .056 25 26 .261 30 27 2.76
GMN .085 49 30 24 .037 .033 24 23 .110 13 13 1.59
DuSEGO-EGNN .067 43 26 22 .026 .031 11 11 .092 11 11 1.54
DuSEGO-SEGNN .072 50 28 26 .031 .033 10 10 .880 10 9 1.62
6.3.3. Experimental Setting.
Due to that SchNet (schutt2017schnet, ), DimeNet++ (gasteiger_dimenetpp_2020, ), SphereNet (liu2021spherical, ), PaiNN (schutt2021equivariant, ), PamNet (zhang2023universal, ) and LEFTNet (du2024new, ) are trained under different train/val/test partitions, we rerun these six baselines, and report the results for a fair comparison. Additionally, we also deploy GMN  (huang2021equivariant, ) and SEGNO  (liu2023improving, ) on the molecule property prediction task and rerun EGNN (satorras2021n, ) and SEGNN (brandstetter2021geometric, ) under the same experimental settings. All these baselines are running on NVIDIA A100 GPUs. The experimental settings are as follows: Adam optimizer with learning rate ${1\times 10^{-3}}$, the number of epochs 1,000, input embedding dim 128, hidden embedding dim 128, output input embedding dim 128, batch size 128, weight decay ${1\times 10^{-12}}$, and layer number 7.
6.3.4. Implementation details.
We build upon the experimental setting introduced in (brandstetter2021geometric, ), where we aim to predict a variety of chemical properties for each of the molecules through regression. We optimized and report the Mean Absolute Error (MAE) between predictions and ground truth. In this section, we employ EGNN and SEGNN as the backbone of DuSEGO, named DuSEGO-EGNN and DuSEGO-SEGNN. In particular, EGNN and SEGNN are implemented based on the implementation details provided in (satorras2021n, ; brandstetter2021geometric, ), which are used as the coupling function for message passing (rusch2022graph, ). Then, we use the ODE solvers provided in torchdiffeq library for solving Eq 8 and Eq. 9.
6.3.5. Results.
Table 5 depicts the overall experimental results of all methods on the QM9 dataset. Compared to the baseline methods, DuSEGO-EGNN achieves the best results for three energy variables ($\Delta\epsilon$, $\epsilon_{\text{HOMO}}$, and $\epsilon_{\text{LUMO}}$), while DuSEGO-SEGNN achieves the best results for four energy variables ($G$, $H$, $U$, and $U_{0}$). Additionally, we observe that incorporating second-order information into EGNN and SEGNN allows DuSEGO-EGNN and DuSEGO-SEGNN to perform better than before in most cases, outperforming GMN and SEGNO in all tasks, which highlights the effectiveness of our approach. As a result, DuSEGO-EGNN and DuSEGO-SEGNN generate competitive results in most property prediction tasks while maintaining relatively simple model structures, without incorporating additional data such as angles and spherical harmonics, as used in PAMNet, SphereNet, PaiNN, and LEFTNet.
7. Conclusion
In this paper, we introduced DuSEGO, a dual second-order equivariant graph neural network designed to address both feature and coordinate representations simultaneously through a graph ODE framework. Our approach not only maintains equivariant properties but also alleviates critical issues in deep graph neural networks, such as over-smoothing and gradient problems. Empirical results on dynamic system modeling, graph autoencoding, and molecular data tasks demonstrate DuSEGO’s superior performance over existing methods, especially in data efficiency and stability across different graph depths and sample sizes. These results underscore DuSEGO’s capacity to model complex physical and molecular systems effectively, even in limited-data scenarios. Future work could explore extending DuSEGO’s application to other domains that benefit from high-order motion laws or developing variations that address specific challenges in large-scale or sparse datasets. Additionally, further integration with third-order or higher-order dynamics could reveal more insights into complex graph-structured data, enhancing its versatility across a wider range of tasks.